What is the solution for this equation? I know i had to use the u, u^2 formula for this.
x^1/2+3x^1/4-10 = 0
First I let u = x^1/4 and u^2 = (x^1/2)^2= 1/2. After that I had the equation
u^2+3u-10=0. I factored it to (u+5)(u-2)=0, and had u=-5,u=2.
(x^1/4)^4= -5^4, (x^1/4)^4= 2^4. My solutions were x=625, x=16. Is this correct?

Answer by Alan3354(69443)

What is the solution for this equation? I know i had to use the u, u^2 formula for this.
x^1/2+3x^1/4-10 = 0
First I let u = x^1/4 and u^2 = (x^1/2)^2= 1/2.
After that I had the equation u^2+3u-10=0.
I factored it to (u+5)(u-2)=0, and had u=-5,u=2.
---------------
(x^1/4)^4= -5^4, (x^1/4)^4= 2^4. My solutions were x=625, x=16. Is this correct?
---------------
They're correct, you can check by subbing them into the original eqn.
For the 625, you have to use -5 4th root of 625 tho, not +5.